OpenStudy (anonymous):

Simplify completely quantity 2 x squared plus 20 x plus 32 all over x squared minus 2 x minus 80.

OpenStudy (anonymous):

A negative 1 over 5 B negative 2 over 5 C 5 x over quantity x minus 5 D 2 open parentheses x plus 2 close parentheses over x minus 10

OpenStudy (anonymous):

\[\frac{2 x^2+20x+32 }{ x^2-2x-80 }\] \[=\frac{ 2\left( x^2+10x+16 \right) }{ x^2-2x-80 }\] 10=8+2 8*2=16 10-8=2 10*-8=-80 make factors and cancel common factor